# Existence for Stable Rotating Star-Planet Systems

### v1

@@ Abstract
This paper investigates the existence and properties of stable, uniformly rotating star-planet systems, i.e. mass ratio is sufficiently small. It is modeled by the Euler-Poisson equations. Following the framework established by McCann for binary stars [46], we adopt a variational approach, and prove the existence of local energy minimizers with respect to the Wasserstein $L^{\infty}$ metric, under the assumed equation of state $P(\rho)=K\rho^{\gamma}$ and under the condition that the mass ratio $m$ is sufficiently small, corresponding to a star–planet system. Such minimizers correspond to solutions of the Euler–Poisson system. We consider two cases. For $\gamma>2$, we not only prove existence but also show, via scaling arguments, that the radii (to be precise, the bounds of the supports of the minimizers) tend to zero. For $\frac{3}{2}<\gamma\leq 2$, we estimate an upper bound for the (potential) expansion rates of the radii, and it turns out that the existence result remains valid in this case as well. Finally, we provide estimates for the distances between different connected components of supports of minimizers and propose a conjecture regarding the number of connected components.
@@ 1 Introduction
In astrophysical fluid dynamics, gas stars and gas planets can be modeled as isolated self-gravitating fluid masses. A fundamental hydrodynamic description for Newtonian stars is provided by the Euler–Poisson system:
@@
Here, $\rho(x,t)\geq 0$ denotes the fluid density at position $x\in\mathbb{R}^{3}$ and time $t\geq 0$, $v(x,t)\in\mathbb{R}^{3}$ is the velocity, and $V(x,t)\in\mathbb{R}$ represents the gravitational potential. The pressure $P(\rho)$ depends solely on density. In this work, we adopt the polytropic equation of state $P(\rho)=K\rho^{\gamma}$.
The motion of stars and planets has long been a classical subject in celestial mechanics, astrophysics, and mathematics. Historical models such as heliocentrism, proposed by Copernicus, and later studies by Newton and others on stellar rotation, form the foundation of rotating star or planet models. Early developments by Maclaurin, Jacobi, Poincaré, and Liapounov focused on incompressible, homogeneous or nearly homogeneous bodies like the Earth. For gaseous bodies such as Jupiter, or systems like galaxies, density inhomogeneity and compressibility need to be accounted for. The first attempts to construct axisymmetric rotating solutions to the compressible Euler–Poisson equations were made by astrophysicists Milne [47] and Zeipel [57] in the early 1920s, followed by Chandrasekhar [11] and Lichtenstein [39] in 1933. Further historical accounts can be found in [12, 33].
Significant progress has been made for the single rotating star problem, where the star rotates about its axis. Two main approaches have been employed to prove the existence of such rotating stars. The first is variational method, as used by Auchmuty and Beals [3, 4] to establish the existence of axisymmetric equilibria for compressible fluids; see also [2, 4, 9, 13, 38, 16]. The second is a perturbative method based on the implicit function theorem, applied near non-rotating Lane–Emden stars [29, 30, 28, 53, 54]. Nonlinear dynamical stability of rotating stars was shown in [43, 44] via variational techniques, while stability theory for non-rotating stars can be found in [18, 32, 49].
These methods are quite general and apply not only to single-body problems but also to $N$-body systems. Variational methods were used by McCann [46] for binary stars and are applied in this paper to star–planet systems. Perturbative approaches appear in the work of Alonso-Orán, Kepka, and Velázquez [1] on incompressible Euler–Poisson systems with an external particle, as well as in Lichtenstein’s construction of rotating binary stars [39].
Each method has its advantages and limitations. For example, the variational approach typically requires $\gamma>\frac{4}{3}$ when $P(\rho)=K\rho^{\gamma}$ to prevent gravitational collapse (see [16, Remark 2.11 and Proposition 2.16] [46, Section 2], or [4, Section 1, 6 and 8]). In contrast, perturbative methods can handle the range $\frac{6}{5}<\gamma<2$ [30], but they rely on prior knowledge of solutions in special non-rotating cases.
When considering two stellar objects, such as binary stars or a star–planet system, we begin with the simplest scenario where individual rotations are neglected, and only orbital revolution is considered. We still refer to such configurations as rotating systems. The construction of rotating binary stars dates back to Lichtenstein [39]. As discussed by Jang and Seok [31], in the $N$-body uniformly rotating case, the Euler–Poisson equations reduce to a single equation:
@@
A notable feature of McCann’s binary stars is that each star admits a variational characterization as a Hamiltonian or energy minimizer under certain constraints, while also being interpretable as a perturbation of simpler objects—such as non-rotating Lane–Emden stars and point-mass relative equilibria (see subsection 5.3 and Section 6 of this paper or [31, Section 3.2]). This characterization is crucial for showing that the support of an energy-constrained minimizer lies in the interior of balls of certain radii. Consequently, the minimizer $\rho$ is in fact a local minimizer in the Wasserstein $L^{\infty}$ distance sense. This ensures that $\nabla P(\rho)$ is well-defined throughout $\mathbb{R}^{3}$, and that (EP’) holds in the whole space $\mathbb{R}^{3}$ (see Theorem 3.1, [46, Theorem 2.1], [15, Section 2]).
@@
Beyond existence, many questions remain for two-body system. For instance, Jang and Seok [31] study the asymptotic profiles, uniqueness, and orbital stability of McCann’s uniformly rotating binary stars. In addition, conjectures concerning the maximal possible number of connected components of the support of solutions are discussed in Section 7 of this paper.
Researchers have also considered systems with more than two stellar objects, such as galaxies or general $N$-star systems. For rotating galaxies, single rotating galaxies were constructed in [52, 53] in the spirit of Lichtenstein and Heilig [28, 39]. There is an extensive literature on non-rotating galaxies: see [6, 7, 24, 50] and references therein. The orbital stability of stationary solutions has seen considerable progress over the past two decades [26, 21, 22, 23, 20, 25, 34, 35, 36, 37, 50, 56], with recent work on linearly oscillating galaxies in [27]. In [10], Campos, Del Pino, and Dolbeault constructed $N$-body rotating galaxies by connecting to relative equilibria in $N$-body dynamics with small uniform angular velocity, perturbing radial equilibria, and solving for the gravitational potential rather than the density, using a finite-dimensional reduction. Although natural, Jang and Seok remark in [31] that the framework in [10] is not ideal for stability analysis. They adapt Rein’s reduction method [49, 50] to extend the discussion of binary star solutions to establish existence and stability for rotating binary galaxies modeled by the Vlasov–Poisson system [31]. Furthermore, they note that for $N\geq 3$, uniformly rotating $N$-body stellar configurations lack a variational structure analogous to the binary case and are not generally expected to be stable [31].
@@
This paper presents the core results extracted from the author’s Master’s thesis [14]. Building upon the preparatory work in two companion papers [15, 16] originating from the same thesis, it investigates the existence and qualitative properties of stable, uniformly rotating star–planet systems governed by the Euler–Poisson equations. This work thereby extends McCann’s framework for binary stars [46] to the setting of small mass ratios, as established in Theorems 2.12 and 2.13.
@@
The paper is structured as follows. Section 2 introduces the variational formulation and states the main results. Section 3 collects preliminary results about two-body systems and non-rotating single star cases from [46, 4, 40, 15, 16]. Section 4, 5 and 6 are the core of this paper, where we will prove existence of star-planet systems. Sections 7 studies distances between (possible) different connected components of planets or stars, leading to the conjecture that minimizer may consist of at most two connected components, corresponding to exactly one star and exactly one planet respectively.
@@ 2.1 Variational formulation
We introduce some notations and variational formulation here, which based on McCann’s construction of binary star solutions, see [46] and [15].
@@
    k{k} dimensional open ball: $B_{R}(a):=\left\{x\in\mathbb{R}^{k}\mid|x-a\mid<R\right\}$ for given $a\in\mathbb{R}^{k},R\in\mathbb{R}$.
@@
    $r(x):=\left(\langle x,x\rangle_{2}\right)^{\frac{1}{2}}=x_{1}^{2}+x_{2}^{2}$.
@@
  8. (viii)
    Given $\delta>0$ the $\delta$-neighbourhood of $\Omega\subset\mathbb{R}^{3}$ to be the set Ω+B_δ(0):=⋃_y ∈Ω{x ∈R^3∣— x-y —¡δ}
  9. (ix)
    $\mu$: Lebesgue measure.
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)$. The fluid interacts with itself through Newtonian gravity hence we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we need to first consider the equation of state for pressure $P(\rho)$. In this paper, we take the polytropic law, i.e.
@@
We first give a “potential admissible class” for $\rho$ and $v$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=1\right\}$ (3)
@@
$U(\rho):=\int A(\rho(x))d^{3}x$ (6)
$G(\sigma,\rho):=\int V_{\sigma}\rho=\iint\frac{\rho(x)\sigma(y)}{|x-y|}d^{3}yd^{3}x$ (7)
$T(\rho,v):=\frac{1}{2}\int|v|^{2}\rho$ (8)
Here $A(\rho)$ is a convex function given in (2), and $U(\rho)$ is internal energy.
@@
$V_{\rho}(x):=\int\frac{\rho(y)}{|y-x|}d^{3}y$ (9)
Hence $G(\rho,\rho)$ is gravitational interaction energy.
We can also see $T(\rho,v)$ is kinetic energy.
@@
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, we have $G(\rho,\rho)<\infty$, see for example [15, Proposition 2.10]. Hence $E(\rho,v)$ is well-defined.
@@
$\bar{x}(\rho):=\frac{\int x\rho(x)d^{3}x}{\int\rho(x)d^{3}x}$ (10)
is at rest. We are interested in finding minimum energy configurations subject to constraints of fixed mass ratio and angular momentum $\boldsymbol{J}$ with respect to the center of mass $\bar{x}(\rho)$. The fluid angular momentum $\boldsymbol{J}(\rho,v)$ is given by:
$\boldsymbol{J}(\rho,v):=\int(x-\bar{x}(\rho))\times v\rho(x)d^{3}x$ (11)
@@
Since the z-component of the angular momentum is specified, the moment of inertia $I(\rho)$ of $\rho$ in the direction of $\hat{e}_{z}$ will be relevant. That is, we define $I(\rho)$ to be moment of inertia of $\rho$ in the direction of $\hat{e}_{z}$:
$I(\rho):=\int\rho r^{2}(x-\bar{x}(\rho))$ (12)
@@
When $\rho$ has positive mass, we have $I(\rho)>0$. See [15, Remark 2.9].
@@
Lemma 2.4 (Expansion of Moment of Inertia [15, Lemma 2.15]).
Let $\rho\geq 0$, $\sigma\geq 0$ be the density functions in $\mathbb{R}^{3}$ with mass $\int\rho=m_{1}<\infty,\int\sigma=m_{2}<\infty$. $\bar{x}(\rho)$ and $\bar{x}(\sigma)$ denote the centers of mass, $I(\rho)$ and $I(\sigma)$ denote the moments of inertia, function $r$ is given in (2.1).
@@
Let support of $\rho$ be the smallest closed set carrying the full mass of $\rho$ (denoted by spt $\rho$). Intuitively we hope spt $\rho$ to be compact as this aligns with the case of a star-planet system and ensures a finite moment of inertia. Therefore, we also introduce a subset $R_{a}\left(\mathbb{R}^{3}\right)$ of $R\left(\mathbb{R}^{3}\right)$:
@@
Since the energy is translation-invariant, we may just consider looking for a energy minimizer $\rho$ such that the center of mass $\bar{x}(\rho)$ is 0, which means $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$.
@@
Given $\rho\in R(\mathbb{R}^{3})$, if we further know $\rho\in L^{\infty}(\mathbb{R}^{3})$, then we know $U(\rho)$ is finite by interpolation inequality (Proposition A.2). This observation, together with Remark 2.2 and Remark 2.3, tells us $E_{J}(\rho)$ is well-defined and finite. For more general pressure assumptions, see [15, Lemma 5.2].
We intend to solve (EP’) by minimizing $E_{J}(\rho)$ and using results from calculus of variations, which involves the variational derivative of energy. We therefore now proceed to define it. We define the perturbation set $P_{\infty}(\rho)$, which depends on energy minimizer $\rho$, as the following:
@@
One can see $P_{\infty}(\rho)$ is a convex cone. And we have the following result:
Lemma 2.6 (Differentiability of Energy $E_{J}(\rho)$ [15, Lemma 5.3, Remark 5.4]).
Given $\rho\in mR(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$, i.e. it is differentiable at $\rho$ in the direction of $P_{\infty}(\rho)$. Moreover, the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{J}^{\prime}(\rho)(\sigma)=\int E_{J}^{\prime}(\rho)\sigma$  [^1] To remain consistent with the notation in [4, Section 4], we use $E_{J}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion.. The function $E_{J}^{\prime}(\rho)$ on the right‑hand side is given by
@@
where $\omega=\frac{J}{I\left(\rho\right)}$ and $R_{\theta}$ is a rotation map about $x_{3}$ axis given in Definition 2.1, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating star-planet system. Here “rotating” means the orbital revolution around each other, while the rotations of objects around their own axes are not considered. Moreover, $E(\rho,v)$ is local minimum, and $({\rho}(t,x),v(t,x))$ is solution to (EP). We will discuss these results in details in Theorem 3.1.
For prescribed angular momentum, we also know the energy $E_{J}(\rho)$ is bounded from below on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by the non-rotating energy $e_{0}$. However, as in Morgan [48], McCann [46, Section 3] demonstrates that this bound — although approached — will not be attained, see also [45]. Therefore, instead of searching for a global energy minimizer in star-planet system, we adopt McCann’s strategy [46]: first establishing the existence of a constrained minimizer of $E_{J}(\rho)$, and then proving that it is also a local minimizer under the appropriate topology induced by Wasserstein $L^{\infty}$ distance.
Due to this uniform rotation observation, in the presence of 2-body systems, although the systems rotate in universe, after fixing a time, we can still assume that they fall within two disjoint regions $\Omega_{m}$ and $\Omega_{1-m}$, widely separated relative to $\frac{J^{2}}{\mu^{2}}$, where $\mu=m(1-m)$ is their reduced mass. Note that the same symbol $\mu$ is also used as for Lebesgue measure, provided no ambiguity arises. For the planet’s mass $m\in(0,1)$ and the star’s mass (1m)(1-m), we consider $E_{J}(\rho)$ is minimized subject to the constraint
$W_{m}:=\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in R\left(\mathbb{R}^{3}\right)\mid\int\rho_{m}=m,\text{spt }\rho_{m}\subset\Omega_{m},\text{spt }\rho_{1-m}\subset\Omega_{1-m}\right\}$ (21)
@@
The reason we set the separation $\eta=\frac{J^{2}}{\mu^{2}}$ in the definitions above is inspired by the Kepler problem. Given two point masses $m$ and 1m1-m, rotating with angular momentum $J>0$ about their fixed center of mass. Assume their separation is $d$, the gravitational energy plus kinetic energy is $\frac{\mu}{d}+\frac{J^{2}}{2\mu d^{2}}$, which reaches its minimum at separation $d=\eta$.
In the star-planet model, when $\eta$ is large, the distance between two objects is large, the gravitational interaction becomes weak. The system thus approximates the ideal case of two isolated, non-rotating, and non-interacting bodies. Therefore, it can be proven that the supports of the objects remain bounded, which sizes are hence negligible compared to the distance between them. It implies the distance between their centers of mass asymptotically approaches that of the point mass model. We will discuss these facts in more details in Section 5 and Section 6, assuming $m$ is sufficiently small, and then establish their connection to the existence of local energy minimizers under topology induced by $W^{\infty}$ (Wasserstein $L^{\infty}$) distance. For the case $J$ is sufficiently large, one can also check McCann’s paper [46, Section 6].
We note the choice of topology for ${R}\left(\mathbb{R}^{3}\right)$ is quite delicate: for $J>0$, it turns out local energy minimizer for $E_{J}(\rho)$ will not exist if the topology of $R(\mathbb{R}^{3})$ is inherited from a topological vector space ([46, Remark 3.7] and [15, Proposition 5.15]). Roughly speaking, the reason is that for any reasonable candidate local minimizer $\rho$, one can always split off a portion of it with very small mass and shift the portion arbitrarily far away, thereby decreasing the energy. However, the resulting configuration still lies in a small neighborhood of $\rho$ with respect to such a topology, which implies that $\rho$ cannot be a local minimizer. See also Remark 3.10.
Hence we need another topology, i.e. the topology induced by the $W^{\infty}$ distance.
@@
$f_{\#}\tilde{\mu}(B):=\tilde{\mu}\left(f^{-1}(B)\right)\quad\text{for any Borel set }B\text{ in }X$
We call $f,g$ the transport maps. Thanks to Strassen’s Theorem, one can check $W^{\infty}$ is truly a metric, as explained in Givens and Shortt [19]. We gives two properties of $W^{\infty}$ distance. For other properties, one can see [46, Lemma 5.1] and [15, Lemma 4.7].
Lemma 2.9 (Simple Properties of the Wasserstein $L^{\infty}$ Metric [46, Lemma 5.1]).

### v2

@@ Abstract
This paper investigates the existence and properties of stable, uniformly rotating star-planet systems, i.e., mass ratio is sufficiently small. It is modeled by the Euler-Poisson equations. Following the framework established by McCann for binary stars [24], we adopt a variational approach, and prove the existence of local energy minimizers with respect to the Wasserstein $L^{\infty}$ metric, under the assumed equation of state $P(\rho)=K\rho^{\gamma}$ and under the condition that the mass ratio $m$ is sufficiently small, corresponding to a star–planet system. Such minimizers correspond to solutions of the Euler–Poisson system. We consider two cases. For $\gamma>2$, we not only prove existence but also show, via scaling arguments, that the radii (to be precise, the bounds of the supports of the minimizers) tend to zero. For $\frac{3}{2}<\gamma\leq 2$, we estimate an upper bound for the (potential) expansion rates of the radii, and it turns out that the existence result remains valid in this case as well. Finally, we provide estimates for the distances between different connected components of supports of minimizers and propose a conjecture regarding the number of connected components.
@@ 1 Introduction
In astrophysical fluid dynamics, gas stars and gas planets can be modeled as isolated self-gravitating fluid masses. A fundamental hydrodynamic description is provided by the Euler–Poisson system:
@@
Here, $\rho(x,t)\geq 0$ denotes the fluid density at position $x\in\mathbb{R}^{3}$ and time $t\geq 0$, $v(x,t)\in\mathbb{R}^{3}$ is the velocity, and $V(x,t)\in\mathbb{R}$ represents the gravitational potential. The pressure $P(\rho)$ depends solely on density. The existence of solutions under different settings has been established via variational methods (e.g., [3, 4, 2, 7, 8, 18, 24]) and perturbative approaches (e.g., [15, 16, 14, 29, 30, 1, 19]).
In this work, we adopt the polytropic equation of state $P(\rho)=K\rho^{\gamma}$, unless otherwise stated. We mainly consider two stellar objects, under the simplifying assumption that individual rotations are neglected and only orbital revolution is considered. We still refer to such configurations as rotating systems. The construction of rotating binary stars dates back to Lichtenstein [19]. As discussed by Jang and Seok [17], in the $N$-body uniformly rotating case, the Euler–Poisson equations reduce to a single equation:
@@
where $\omega>0$ is the angular velocity, $\tilde{\rho}$ is a compactly supported density function, and $V_{\tilde{\rho}}(x)={\int_{\mathbb{R}^{3}}{\frac{\tilde{\rho}(y)}{\left|{y-x}\right|}\,dy}}$. As shown in [10, Section 2], if $\tilde{\rho}$ satisfies the reduced Euler–Poisson equation (EP’), then $\rho\left({t,y}\right)=\tilde{\rho}\left(R_{-\omega t}y\right)$ and $v\left({t,y}\right)=\omega\left({-y_{2},y_{1},0}\right)^{T}$ satisfy the original system (EP), where $R_{-\omega t}y$ is the rotation map given in Definition 2.1. In this case, it is easy to see that $\Delta v=0$, so $(\rho,v)$ also satisfies the Navier-Stokes-Poisson system. Moreover, since $\nabla\cdot v=0$, no compression occurs in this uniform rotation scenario. Nevertheless, based on physical intuition, we still regard this as a compressible model, corresponding to gaseous stars and planets.
@@
A notable feature of McCann’s framework is that each star admits a variational characterization as a Hamiltonian or energy minimizer under certain constraints, while also being interpretable as a perturbation of simpler objects—such as non-rotating Lane–Emden stars and point-mass relative equilibria (see subsection 5.3 and Section 6 of this paper or [17, Section 3.2]). This characterization is crucial for showing that the support of an energy-constrained minimizer lies in the interior of balls of certain radii. Consequently, the minimizer $\rho$ is in fact a local minimizer in the Wasserstein $L^{\infty}$ distance sense. This ensures that $\nabla P(\rho)$ is well-defined throughout $\mathbb{R}^{3}$, and that (EP’) holds in the whole space $\mathbb{R}^{3}$ (see Theorem 3.1, [24, Theorem 2.1], [10, Section 2]).
@@
Beyond existence, many questions remain for two-body systems. For instance, conjectures concerning the maximal possible number of connected components of the support of solution are discussed in Section 7 of this paper. Jang and Seok [17] study the asymptotic profiles, uniqueness, and orbital stability of McCann’s uniformly rotating binary stars. They also establish existence and stability for rotating binary galaxies modeled by the Vlasov–Poisson system [17], a work that extends the discussion of binary star solutions by adapting Rein’s reduction method [26, 27].
@@
This paper presents the core results extracted from the author’s Master’s thesis [9]. It investigates the existence and qualitative properties of stable, uniformly rotating star–planet systems governed by the Euler–Poisson equations, as established in Theorems 2.12 and 2.13.
Although this paper builds upon the preparatory work in two companion papers [10, 11] originating from the same thesis [9], it is largely self-contained. Relevant auxiliary results can also be found in other existing literature cited at the corresponding locations herein, with the author’s papers [10, 11] providing refined and more complete proofs.
@@
The paper is structured as follows. Section 2 introduces the variational formulation and states the main results. Section 3 collects preliminary results about two-body systems and non-rotating single star cases from [24, 4, 20, 10, 11]. Sections 4, 5 and 6 are the core of this paper, where we will prove existence of star-planet systems. Sections 7 studies distances between (possible) different connected components of planets or stars, leading to the conjecture that minimizer may consist of at most two connected components, corresponding to exactly one star and exactly one planet respectively.
@@ 2.1 Variational formulation
We introduce some notations and variational formulation here, which are based on McCann’s construction of binary star solutions, see [24] and [10].
@@
    $r(x):=\left(\langle x,x\rangle_{2}\right)^{\frac{1}{2}}=\sqrt{x_{1}^{2}+x_{2}^{2}}$.
@@
    Rotation map: R_θ:=(cosθ-sinθ0sinθcosθ0001)
@@
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)$. The fluid interacts with itself through Newtonian gravity hence we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we need to first consider the equation of state for pressure $P(\rho)$. In this paper, we take the polytropic law, i.e.,
@@
We first give “admissible classes” for $\rho$ and $v$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=1\right\}$ (3)
@@
$U(\rho):=\int_{\mathbb{R}^{3}}A(\rho(x))\,dx$ (6)
$G(\sigma,\rho):=\int_{\mathbb{R}^{3}}V_{\sigma}\rho\,dx=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\rho(x)\sigma(y)}{|x-y|}\,dy\,dx$ (7)
$T(\rho,v):=\frac{1}{2}\int_{\mathbb{R}^{3}}|v|^{2}\rho\,dx$ (8)
Here $A(\rho)$ is a convex function given in (2), and $U(\rho)$ is the internal energy.
@@
$V_{\rho}(x):=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|y-x|}\,dy$ (9)
Hence $G(\rho,\rho)$ is the gravitational potential energy (also called gravitational interaction energy), and $T(\rho,v)$ is the kinetic energy.
@@
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, we have $G(\rho,\rho)<\infty$, see for example [10, Section 2]. Hence $E(\rho,v)$ is well-defined.
@@
$\bar{x}(\rho):=\frac{\int_{\mathbb{R}^{3}}x\rho(x)\,dx}{\int_{\mathbb{R}^{3}}\rho(x)\,dx}$ (10)
is at rest. We are interested in finding minimum energy configurations subject to constraints of fixed mass ratio and fixed angular momentum $\boldsymbol{J}$ with respect to the center of mass $\bar{x}(\rho)$. The fluid angular momentum $\boldsymbol{J}(\rho,v)$ is given by:
$\boldsymbol{J}(\rho,v):=\int_{\mathbb{R}^{3}}(x-\bar{x}(\rho))\times v\rho(x)\,dx$ (11)
@@
Since the z-component of the angular momentum is specified, the moment of inertia $I(\rho)$ of $\rho$ in the direction of $\hat{e}_{z}$ will be relevant. That is, we define the moment of inertia of $\rho$ in the direction of $\hat{e}_{z}$, denoted by $I(\rho)$, as follows:
$I(\rho):=\int_{\mathbb{R}^{3}}\rho r^{2}(x-\bar{x}(\rho))\,dx=\int_{\mathbb{R}^{3}}\rho(x)\left((x_{1}-\bar{x}(\rho)_{1})^{2}+(x_{2}-\bar{x}(\rho)_{2})^{2}\right)\,dx$ (12)
@@
When $\rho$ has positive mass, we have $I(\rho)>0$. See [10, Section 2].
@@
Lemma 2.4 (Expansion of Moment of Inertia [10, Section 2]).
Let $\rho\geq 0$, $\sigma\geq 0$ be the density functions in $\mathbb{R}^{3}$ with mass $\int_{\mathbb{R}^{3}}\rho\,dx=m_{1}<\infty$, $\int_{\mathbb{R}^{3}}\sigma\,dx=m_{2}<\infty$. $\bar{x}(\rho)$ and $\bar{x}(\sigma)$ denote the centers of mass, $I(\rho)$ and $I(\sigma)$ denote the moments of inertia, function $r$ is given in (2.1).
@@
Let the support of $\rho$ be the smallest closed set carrying the full mass of $\rho$ (denoted by spt $\rho$). Intuitively we hope spt $\rho$ to be compact as this aligns with the case of a star-planet system and ensures a finite moment of inertia. Therefore, we also introduce a subset $R_{a}\left(\mathbb{R}^{3}\right)$ of $R\left(\mathbb{R}^{3}\right)$:
@@
Since the energy is translation-invariant, we may just consider looking for an energy minimizer $\rho$ such that the center of mass $\bar{x}(\rho)$ is 0, which means $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$.
@@
Given $\rho\in R(\mathbb{R}^{3})$, if we further know $\rho\in L^{\infty}(\mathbb{R}^{3})$, then we know $U(\rho)$ is finite by interpolation inequality [6, Section 4.2]:
$\|\rho\|_{L^{r}}\leq\|\rho\|_{L^{p}}^{\theta}\|\rho\|_{L^{q}}^{1-\theta},\quad\text{for }p\leq r\leq q,$ (17)
where $\frac{1}{r}=\frac{\theta}{p}+\frac{1-\theta}{q}$ for some $\theta\in[0,1]$. This observation, together with Remark 2.2 and Remark 2.3, tells us $E_{J}(\rho)$ is well-defined and finite. Similar results hold for more general pressures; see [10, Section 5].
We intend to solve (EP’) by minimizing $E_{J}(\rho)$ and using results from calculus of variations, which involves the variational derivative of energy. We therefore now proceed to define it. We first define the perturbation set $P_{\infty}(\rho)$, which depends on energy minimizer $\rho$, as the following:
@@
One can see $P_{\infty}(\rho)$ is a convex cone. And then we have the following result:
Lemma 2.6 (Differentiability of Energy $E_{J}(\rho)$ [10, Section 5]).
Given $\rho\in mR(\mathbb{R}^{3})$ with $U(\rho)<\infty$, $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$, i.e. it is differentiable at $\rho$ in the direction of $P_{\infty}(\rho)$. Moreover, the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{J}^{\prime}(\rho)(\sigma)=\int_{\mathbb{R}^{3}}E_{J}^{\prime}(\rho)\sigma\,dx$  [^1] To remain consistent with the notation in [4, Section 4], we use $E_{J}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion.. The function $E_{J}^{\prime}(\rho)$ on the right‑hand side is given by
@@
where $\omega=\frac{J}{I\left(\rho\right)}$ and $R_{\theta}$ is a rotation map about $x_{3}$ axis given in Definition 2.1, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating star-planet system. Here “rotating” means the orbital revolution around each other, while the rotations of objects around their own axes are not considered. Moreover, $E(\rho,v)$ is local minimum, and $({\rho}(t,x),v(t,x))$ is solution to (EP). We will discuss these results in detail in Theorem 3.1.
For prescribed angular momentum, we also know the energy $E_{J}(\rho)$ is bounded from below on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by the non-rotating energy $e_{0}$. However, as in Morgan [25], McCann [24, Section 3] demonstrates that this bound — although approached — will not be attained, see also [23, 10]. Therefore, instead of searching for a global energy minimizer in star-planet system, we adopt McCann’s strategy [24]: first establishing the existence of a constrained minimizer of $E_{J}(\rho)$, and then proving that it is also a local minimizer under the appropriate topology induced by Wasserstein $L^{\infty}$ distance.
Due to this uniform rotation observation, in the presence of 2-body systems, although the systems rotate in universe, after fixing a time, we can still assume that they fall within two disjoint regions $\Omega_{m}$ and $\Omega_{1-m}$, widely separated relative to $\frac{J^{2}}{\mu_{r}^{2}}$, where $\mu_{r}=m(1-m)$ is their reduced mass. For the planet’s mass $m\in(0,1)$ and the star’s mass (1m)(1-m), we consider $E_{J}(\rho)$ is minimized subject to the constraint
$W_{m}:=\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in R\left(\mathbb{R}^{3}\right)\mid\int_{\mathbb{R}^{3}}\rho_{m}\,dx=m,\text{spt }\rho_{m}\subset\Omega_{m},\text{spt }\rho_{1-m}\subset\Omega_{1-m}\right\}$ (22)
@@
The reason we set the separation $\eta=\frac{J^{2}}{\mu_{r}^{2}}$ in the definitions above is inspired by the Kepler problem. Given two point masses $m$ and 1m1-m, rotating with angular momentum $J>0$ about their fixed center of mass, if we assume their separation is $d$, then the gravitational energy plus kinetic energy is $\frac{\mu_{r}}{d}+\frac{J^{2}}{2\mu_{r}d^{2}}$, which reaches its minimum at separation $d=\eta$.
In the star-planet model, when $\eta$ is large, the distance between two objects is large, and the gravitational interaction becomes weak. The system thus approximates the ideal case of two isolated, non-rotating, and non-interacting bodies. Therefore, it can be proven that the supports of the objects remain bounded, whose sizes are hence negligible compared to the distance between them. It implies the distance between their centers of mass asymptotically approaches that of the point mass model. We will discuss these facts in more detail in Section 5 and Section 6, assuming $m$ is sufficiently small. We will then establish their connection to the existence of local energy minimizers under the topology induced by $W^{\infty}$ (Wasserstein $L^{\infty}$) distance. For the case $J$ is sufficiently large, one can also check McCann’s paper [24, Section 6].
We note the choice of topology for ${R}\left(\mathbb{R}^{3}\right)$ is quite delicate: for $J>0$, it turns out local energy minimizer for $E_{J}(\rho)$ will not exist if the topology of $R(\mathbb{R}^{3})$ is inherited from a topological vector space ([24, Remark 3.7] and [10, Section 5]). Roughly speaking, the reason is that for any reasonable candidate local minimizer $\rho$, one can always split off a portion of it with very small mass and shift the portion arbitrarily far away, thereby decreasing the energy. However, the resulting configuration still lies in a small neighborhood of $\rho$ with respect to such topology, which implies that $\rho$ cannot be a local minimizer.
Hence we need another topology, i.e., the topology induced by the $W^{\infty}$ distance.
@@
[…]
We call $f,g$ the transport maps. Thanks to Strassen’s Theorem, one can check $W^{\infty}$ is truly a metric, as explained in Givens and Shortt [13]. We give two properties of the $W^{\infty}$ distance. For other properties, one can see [24, Lemma 5.1] and [10, Section 4].
Lemma 2.9 (Simple Properties of the Wasserstein $L^{\infty}$ Metric [24, Lemma 5.1]).